Please help me solve this equation
Use the echelon method to solve this system of two equations in two unknowns
x/2 + y = 3/2
x/3 + y = 1/3

Answer by rothauserc(4718)

x/2 + y = 3/2
x/3 + y = 1/3
subtract second equation from first
x/6 = 7/6
multiply both sides of = by 6
x = 7
substitute for x in first equation
7/2 + y = 3/2
y = -4/2 = -2
solution is x = 7, y = -2
check answer by substituting in second equation
7/3 -2 = 1/3
1/3 = 1/3
our answer checks :-)